the kth term of the series is given by 2k3k the sum of first 100 terms ?

Grade:12th pass

1 Answers

Arun
5 years ago
Dear Hamid
 
kth term Tk = 6^k
 
hence
 
Sum = Sk = 6^1 + 6^2 + ….. + 6^k
 
 Hence sum of 100 terms = S = 6 *(6^100 – 1) / (6 – 1)
 = (6^101 – 6 )/ 5